# Motions of spinning particles and chaos bound in Reissner-Nordström spacetime

### v2

@@ 4 Conclusions and discussions
In this work, we investigated the motion of the spinning particles around the RN BH, and discussed the relationship between the LEs of their classical chaotic motion and the chaos bound. Taking into account the two cases where the spin direction of the particle is aligned and anti-aligned with its total angular momentum direction, we found that the exponents can surpass the bound under the specific conditions. Therefore, the phenomenon of the chaos bound violation is also observable within the spinor field.
@@
When the particle is charged, we took into account the influence of the electromagnetic force. It was observed that there are certain similarities in the variation of the exponent with the spin between the charged and neutral particles (except for the case where the angular momentum is $5.00$). This indicates that the electromagnetic force cannot alter the trend of how the exponent varies with respect to the spin and angular momentum, but it can modify the value of the exponent. The likely reason for this phenomenon is that the electromagnetic force is not sufficiently strong. If both the BH’s and particle’s charges are large enough simultaneously, the trend of the exponent’s variation may change.
There are two explanations for the violation of the bound. The first explanation asserts that there is no real violation. Instead, it requires a modification HT5 ; LTW1 . The second explanation admits the violation, and its proponents posit that this violation may be associated with the thermodynamic stability of BHs. However, the subsequent research has demonstrated that when the BH undergoes the phase transition, the violation occurs in the BH branch within its stable region LGD .

### v3

@@ 4 Conclusions and discussions
In this work, we explored the motions of the spinning particles around the RN BH, and examined the chaos bound. Taking into account the two cases where the spin direction of the particle is aligned and anti-aligned with its total angular momentum direction, we found that the LEs can surpass the surface gravity of this BH under the specific conditions. Therefore, the phenomenon of the chaos bound violation was found in the spinor field.
@@
When the particle is charged, we took into account the influence of the electromagnetic force. It was observed that there are certain similarities in the variation of the exponent with the spin between the charged and neutral particles (except for the case where the angular momentum is $5.00$). This indicates that the electromagnetic force cannot change the trend of how the exponent varies with the spin and angular momentum. However, it can alter the values of the exponent and result in the violation. The likely reason for this phenomenon is that the electromagnetic force is not sufficiently strong. If both the BH’s and particle’s charges are large enough simultaneously, the trend of the exponent’s variation may change.
During the discussion, since factors such as the backreaction of the particles on the background spacetime have not been taken into consideration, the values of the exponents will change once these factors are considered. There are two explanations for the violation. The first explanation asserts that there is no real violation, and thus the bound requires modification HT5 ; LTW1 . The second explanation admits the violation, and its proponents posit that this violation may be associated with the thermodynamic stability of BHs. However, the subsequent research has demonstrated that when the BH undergoes the phase transition, the violation occurs in the BH branch within its stable region LGD .